OpenStudy (anonymous):

please help! find the vertex, focus,and directrix of the parabola and sketch its graph (x+3/2)^2=4(y-2)

OpenStudy (experimentx):

rewrite it as \[ y = \frac 14 (x+3/2)^2 + 2\] put, x=-3/2, y = 2 , this is your vertex

OpenStudy (experimentx):

since the parabola is a vertical parabola, the focus is going to be (-3/2,a)  let directrix be y=b